OpenStudy (anonymous):

What number is midway between 2/5 and 1?

OpenStudy (anonymous):

I really suck on these O_o

OpenStudy (anonymous):

use the formula: \[\frac{ a+b}{2 }\] a= 2/5 b= 1

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

okay so you do 2/5+1 divided by 2 so it equals 7/10?

OpenStudy (anonymous):

so the answer is 7/10?

OpenStudy (skullpatrol):

Yes

OpenStudy (anonymous):

Thanks B)

OpenStudy (skullpatrol):

Thanks for trying :)

OpenStudy (anonymous):

@Algebra? yes.. Is it clear?

OpenStudy (anonymous):

yessss haha got it thank you ^-^

OpenStudy (anonymous):

welcome